tajah33:

A caterer expects a minimum of 160 but no more than 192 people at a wedding. One fruit tray serves 16 people. The compound inequality represents x, the number of fruit trays the caterer should bring to the wedding. 160 ≤ 16x≤ 192 Which is a possible number of fruit trays the caterer might bring to the wedding?

Hero:

@tajah33 Hint: Divide each side by 16 to find x: \(\dfrac{160}{16} \le \dfrac{16x}{16} \le \dfrac{192}{16}\)